a capacitor C=0.01F in series with a resistor R=20 ohm is charged from a batter E=10V . assuming that initially the capacitor is completely uncharged , determine the charge Q(t) and the current i(t) at any time t.


1
Expert's answer
2021-11-22T06:31:12-0500
"RQ'+\dfrac{1}{C}Q=V"

"20Q'+\dfrac{1}{0.01}Q=10"

"Q'+5Q=0.5"

"Q'=-5(Q-0.1)"

"\dfrac{dQ}{Q-0.1}=-5dt"

Integrate


"Q=0.1+c_1e^{-5t}"

Given "Q(0)=0"

"0=0.1+c_1"

"c_1=-0.1"

a)

"Q(t)=0.1-0.1e^{-5t}"

b)


"i(t)=Q'(t)=0.5e^{-5t}"